OpenStudy (jchick):

Choose the slope-intercept equation of the line that passes through the point (2, -3) and is perpendicular to y = 2/3x + 3

OpenStudy (jchick):

a. y = -three halvesx b. y = three halvesx - 6 c. y = -two thirdsx - 5 thirds d. y = two thirdsx - 13 thirds

OpenStudy (solomonzelman):

perpendicular to the line y = 2/3x + 3, means with a slope of -3/2.

OpenStudy (solomonzelman):

So you have the point (2,-3) mand the slope (-3/2).

OpenStudy (solomonzelman):

\[y-y_1=\left(x-x_1\right)\]

OpenStudy (misty1212):

lol good thing mathematicians invented fraction notation so you can write \(\frac{3}{2}\) etc

OpenStudy (solomonzelman):

I don't have my latex thing with me right now

OpenStudy (jchick):

A

OpenStudy (solomonzelman):

yes

OpenStudy (jchick):

Ok thanks I get it now

OpenStudy (solomonzelman):

:)